# Confidence-weighted integration of human and machine judgments for superior decision-making

### v1

@@ 2.6 Cross-validation procedure
Given the 503 observations for the 100 test cases, we performed a leave-one-out cross-validation (LOOCV) [17, 18] that provides the best bias-variance trade-off for small datasets. Consider the evaluation of the $i$-th test case. With this procedure, we removed all instances of test case $i$ (between 2 and 9) leaving the remaining 99 test cases with all their instances (between 494 and 501) to train the classifier teams. For testing, we utilized all the instances of test case $i$. This was repeated for all 100 test cases, yielding 503 predictions. Note that for individual teammates the evaluation comprised only the testing phase. When the predicted labels of any team or individual teammate in this study were randomly shuffled, the LOOCV accuracy dropped approximately to chance level (i.e., 50%).
3 Results
Figure 2: Conditions for effective collaboration between human experts and LLMs were satisfied. a) When LLMs were confident in their BrainBench judgments, they were more likely to be correct. Confidence ratings were sorted into equal bins, and the mean accuracy for each bin was plotted. The positive slope of the black regression lines for Llama2 chat models (7B, 13B, and 70B) indicates well-calibrated confidence, meaning higher confidence correlated with higher accuracy. b) LLMs surpassed human experts on BrainBench overall. Error bars represent standard error of the mean using a binomial model. c) Item difficulty Spearman correlations among LLMs and human experts. For LLMs, difference in perplexity between incorrect and correct abstracts was used to determine the relative difficulty of test cases. Mean accuracy was used for human experts. LLMs align more with each other than humans, which implies human-machine teams will be diverse.
Figure 3: Performance of all possible teams using the confidence-weighted logistic combination model. Adding a human to a team with one or more machines (blue points) always has a benefit (green points). Llama2 chat 7B, 13B, and 70B models were considered. Each data point corresponds to the average across 503 evaluations. Error bars represent standard error of the mean using a binomial model.
We first assessed the conditions for effective collaboration (i.e., complementarity), namely calibration of confidence and classification diversity among team members. Both humans and LLMs were calibrated in that accuracy was positively correlated with confidence (Figure 2a). Diversity held in that LLMs and humans differed on which test items led to errors (Figure 2c).
In terms of accuracy (Figure 2b), LLMs numerically surpassed humans by a small margin ($t(2)=5.20$, $P<0.05$). Thus, we can consider whether humans can benefit teams consisting of machines that perform comparably or better.
Of primary interest was whether teams including humans performed better than LLM only teams. All 15 possible team combinations, ranging from individual teammates to a 4-way human-LLM team, were considered (Figure 3). Adding a human teammate to LLM-only teams always improved the team’s performance ($t(6)=8.22$, $P<0.0001$).
Our confidence-weighted logistic regression approach follows from the principles of a Bayesian combination model that fosters human-machine complementarity [12]. One questions is how well our logistic regresion approach compares to the Bayesian approach. Our confidence-weighted regression model compared favorably (Figure 4) to the Bayesian model when evaluated on the three human-LLM teams for which the Bayesian model is intended to apply ($t(2)=4.08$, $P<0.01$). This success is impressive given that the regression approach takes seconds to compute on a current desktop whereas the Bayesian approach would take 12-13 days.
Figure 4: Comparison between Bayesian and confidence-weighted logistic combination models for human-LLM teams. Each data point corresponds to the average across 503 evaluations. Llama2 chat 7B, 13B, and 70B models were considered. Error bars represent standard error of the mean using a binomial model. The confidence-weighted logistic combination model more effectively integrated human and machine judgments.
4 Discussion
Can humans team effectively with LLMs when the humans perform worse? We developed a confidence-weighted regression approach that can integrate judgments from any number of teammates. Using this method and testing on a forecasting benchmark (Figure 1), we found that human-LLM teams achieve complementarity, that is their combined performance bests that of either teammate alone (Figure 3). Complementarity was achieved because two critical conditions were satisfied, namely confidence was well-calibrated and classification diversity held among teammates (Figure 2). Strikingly, every combination of LLMs benefited from adding a human to the team (Figure 3).
Our approach was informed by a Bayesian method for combining judgments of humans and machines [12]. Our approach has a number of advantages, including ease of implementation, very fast runtime, an interpretable solution, and readily extendable to any number of teammates. Surprisingly, our confidence-weighted regression approach performed better than the Bayesian approach (Figure 4). One possibility is that the discretization of continuous confidence measures, which the Bayesian model requires, limited its performance. Perhaps an alternative formulation would perform better. Unfortunately, reformulating the Bayesian model and properly implementing it requires substantial effort and expertise. In contrast, because our confidence-weighted integration model is formulated within a regression framework, it is straightforward to extend to include other factors. For example, nonlinear relationships (e.g., polynomial terms) between confidence-weighted predicitons and outcomes could be considered.
This study explored the possibility of a collaborative approach between humans and LLMs for superior decision-making in predicting neuroscience outcomes. Our confidence-weighted regression method effectively combined human and LLM judgments because teammates fulfilled the conditions of well-calibrated confidence and classification diversity. Our results suggest there is a place for humans in teams with machines even when the machines perform better. We hope our work facilitates successful collaborations between humans and machines in addressing important challenges.
5 Data availability
The human participant data and LLM perplexity scores utilized in this study have been previously reported in [1], and are publicly available at
https://github.com/braingpt-lovelab/BrainBench .
6 Code availability
All computer code associated with this work including combination model implementations, team evaluations, and analyses are publicly available at
https://github.com/braingpt-lovelab/haico .
@@ Acknowledgments
This work was supported the ESRC (ES/W007347/1), Microsoft (Accelerate Foundation Models Research Program), and a Royal Society Wolfson Fellowship (18302) to B.C.L.
@@
Algorithm 1 Bayesian Combination Model
@@
Algorithm 1 illustrates the inference procedure of the Bayesian combination model, which comprises two stages: estimation of parameters (training) and prediction of label probabilities (testing). For the estimation of parameters, the given data is: true labels $z$, probability confidence scores $\pi_{M}$, and human classification $y$ and confidence ratings $r$. We assumed that all human participants shared the same set of parameters ($a_{H}$ , $b_{H}$ , $\sigma_{H}$, $c$, $\delta$, and $\tau$). Human confidence ratings on the slider bar were mapped to range between 1 and 100. A wide range is computationally expensive, thus, we aggregate it into three levels: “0: low confidence”, “1: moderate confidence”, and “2: high confidence”. Then, the aggregated confidence rating used for analysis, $r\in\{0,1,2\}^{N}$, reads
@@
The utilized cutpoints (i.e., 33 and 66) produced a good agreement between confidence and accuracy. Among the evaluations of human participants, “low” had $63.2\%$ average accuracy ($n=174$), “moderate” had $66.5\%$ ($n=185$), and “high” had $81.9\%$ ($n=144$). To estimate the posterior over the underlying parameters, a No-U-Turn Sampler (NUTS) for Markov chain Monte Carlo (MCMC) [19] was used with $n_{w}=500$ warmup steps, $n_{c}=3$ chains, and $n_{s}=25$ samples. For the prediction of label probabilities the given data is the same as before but $z$, that is now latent. The prior distributions were set according to [12]. The correlation between team members was $\rho\sim\mathrm{Uniform}(-1,1)$. For the LLM classifier, $a\sim\mathcal{N}(0,10)$, $b\sim\mathcal{N}(0,10)$, and $\sigma\sim\mathrm{Uniform}(0,15)$. For human classifications, $a$ was the same but $b$ and $\sigma$ were adjusted to $b=0$ and $\sigma=1$ for the purpose of identifiability [12]. In addition, we used $\delta\sim\mathrm{Uniform}(0,100)$ for the scaling parameter and uniform priors on the cutpoints, $c\sim\mathrm{Uniform}(0,1)$, with the constraint that the cutpoints are ordered (i.e. $c_{i}<c_{i+1}\;\forall\,i=1,\ldots,R-2)$. Finally, we set $\tau=0.05$ for best convergence results [12].

### v2

@@ 4 Methods
Given $M$ observations for $N$ test cases (with $M>N$), we performed a leave-one-out cross-validation (LOOCV) [25, 26] that provides the best bias-variance trade-off for small datasets. Consider the evaluation of the $i$-th test case. With this procedure, we removed all instances of test case $i$ leaving the remaining $N-1$ test cases with all their instances to train the classifier teams. For testing, we utilized all the instances of test case $i$. This was repeated for all $N$ test cases, yielding $M$ predictions. Note that for individual teammates the evaluation comprised only the testing phase. When the predicted labels of any team or individual teammate in this study were randomly shuffled, the LOOCV accuracy dropped approximately to chance level (i.e., $1/L$).
@@ Acknowledgments
This work was supported by a NeuroData Discovery Award from the Kavli Foundation to F.Y., the ESRC (ES/W007347/1), Microsoft (Accelerate Foundation Models Research Program), a Royal Society Wolfson Fellowship (18302), and an AI safety grant from the Foresight Institute to B.C.L.
Author contributions
Conceptualization, F.Y. and B.C.L; methodology, F.Y., X.L., and B.C.L; investigation, F.Y. and X.L.; writing-–original draft, F.Y.; writing-–review & editing, F.Y., X.L., and B.C.L; funding acquisition, F.Y. and B.C.L; resources, F.Y. and O.V.M.; supervision, B.C.L.
Declaration of interests
The authors declare no competing interests.
@@
Algorithm S1 Bayesian Combination Model
@@
Algorithm S1 illustrates the inference procedure of the Bayesian combination model, which comprises two stages: estimation of parameters (training) and prediction of label probabilities (testing). For the estimation of parameters, the given data is: true labels $z$, probability confidence scores $\pi_{M}$, and human classification $y$ and confidence ratings $r$. We assumed that all human participants shared the same set of parameters ($a_{H}$ , $b_{H}$ , $\sigma_{H}$, $c$, $\delta$, and $\tau$). In the neuroscience forecasting task, human confidence ratings on the slider bar were mapped to range between 1 and 100. A wide range is computationally expensive, thus, we aggregate it into three levels: “0: low confidence”, “1: moderate confidence”, and “2: high confidence”. Then, the aggregated confidence rating used for analysis, $r\in\{0,1,2\}^{N}$, reads
@@
The utilized cutpoints (i.e., 33 and 66) produced a good agreement between confidence and accuracy. Among the evaluations of human participants, “low” had $63.2\%$ average accuracy ($n=174$), “moderate” had $66.5\%$ ($n=185$), and “high” had $81.9\%$ ($n=144$).
To estimate the posterior over the underlying parameters, a No-U-Turn Sampler (NUTS) for Markov chain Monte Carlo (MCMC) [27] was used with $n_{w}=500$ warmup steps, $n_{c}=3$ chains, and $n_{s}=25$ samples. For the prediction of label probabilities the given data is the same as before but $z$, that is now latent. The prior distributions were set according to [13]. The correlation between team members was $\rho\sim\mathrm{Uniform}(-1,1)$. For the LLM classifier, $a\sim\mathcal{N}(0,10)$, $b\sim\mathcal{N}(0,10)$, and $\sigma\sim\mathrm{Uniform}(0,15)$. For human classifications, $a$ was the same but $b$ and $\sigma$ were adjusted to $b=0$ and $\sigma=1$ for the purpose of identifiability [13]. In addition, we used $\delta\sim\mathrm{Uniform}(0,100)$ for the scaling parameter and uniform priors on the cutpoints, $c\sim\mathrm{Uniform}(0,1)$, with the constraint that the cutpoints are ordered (i.e. $c_{i}<c_{i+1}\;\forall\,i=1,\ldots,R-2)$. Finally, we set $\tau=0.05$ for best convergence results [13].
Figure S1: Performance of the Bayesian and confidence-weighted logistic combination models in a low noise object recognition task [13]. Accuracy results on low levels of image noise ($\Omega=80$) with the Bayesian and confidence-weighted combination models, where humans outperform machines ($t(4)=-4.77$, $P<0.01$). Each data point corresponds to the average across 7 247 image evaluations. Error bars represent standard error of the mean using a binomial model. The performance between the Bayesian and confidence-weighted combination models is indistinguishable for human-machine teams ($t_{Welch}(6)=0.49$, $P=0.32$). However, the confidence-weighted combination model presented improved results for machine-machine teams ($t_{Welch}(16)=3.75$, $P<0.001$).
Figure S2: Removing confidence from the confidence-weighted logistic combination model does not negatively impact team performance in a noisy object recognition task [13]. Accuracy results on high levels of image noise ($\Omega=125$) with the confidence-weighted logistic combination model, where confidence scores were set to 1, i.e., $f(x)=1$ in Eq. (3). Each data point corresponds to the average across 7 239 image evaluations. Error bars represent standard error of the mean using a binomial model. Similarly as in the standard scenario (Figure 1), human-machine teams surpass machine-only teams ($t_{Welch}(24)=2.80$, $P<0.01$), although with a smaller margin.
Figure S3: Optimal squashing in the confidence-weighted logistic combination model does not improve overall team performance in the neuroscience forecasting task [14]. Accuracy results on the neuroscience forecasting task with the confidence-weighted logistic combination model, where the parameter $\alpha$ in Eq. (3) is optimized. Llama2 chat 7B, 13B, and 70B models were considered. Each data point corresponds to the average across 503 test case evaluations. Error bars represent standard error of the mean using a binomial model. Similarly as in the standard scenario (Figure 4), human-LLM teams surpass LLM-only teams ($t_{Welch}(11)=7.20$, $P<0.0001$). The performance between standard and squashing scenarios is, however, indistinguishable ($t_{Welch}(19)=0.05$, $P=0.48$).
Figure S4: Adding interaction terms to the confidence-weighted logistic combination model does not improve overall team performance in the neuroscience forecasting task [14]. Accuracy results of the confidence-weighted logistic combination model on the neuroscience forecasting task as a function of the degree of the polynomial expansion of the features. Llama2 chat 7B, 13B, and 70B models were considered. Each data point corresponds to the average across 503 test case evaluations. Only interactions were considered, as the results including pure terms is identical. Team performance is independent of interaction terms.